OpenStudy (anonymous):

Open Solution Problem;Soft snow can be compressed by pressure about 3.0 ×10^3 Pa. What is the smallest area that a pair of snowshoes must have if they will enable a 80 kg person to walk over the snow without sinking in? Consider g = 9.8 m/s2.

OpenStudy (anonymous):

The first person who gets it earns a medal and fan

OpenStudy (anonymous):

First you have to consider the ?Newtonian force acting on the ground. This is caculated using Mg

OpenStudy (anonymous):

And use the equation Pressure=F/A(surface area in m^2) with inequality equation

OpenStudy (ddcamp):

First find the force the person will exert on the snow, using Newton's 2nd law: \[F = ma \\ F = 88kg * 9.8 \frac{m}{s^2}\]

OpenStudy (anonymous):

Correct! Newton's second law or simply mass*gravityt

OpenStudy (anonymous):

In fact you are calculating the normal force using that equation above

OpenStudy (ddcamp):

Use the equation for pressure:\[P = F/A \\ 3000 Pa = 862.4 N / A \\ A = 862.4N / 3000 Pa \\ A = 0.287 m^2\]

OpenStudy (anonymous):

Well done.